oneday 45 frogs were captured from a pond, marked and then returned to the pond.
The next day, 40 frogs were observed in the pond,of which 10 had been marked the previous day. Assuming the marked frogs are equally distributed among all the frogs in the pond, what is the best estimate of how many frogs live in the pond?

Answer by khwang(438)

Assume totally, there are x frogs.
ratio of being amrked:
45 / x = 10/40 = 1/4
so x = 180 (if no dying or new born in one day)
Kenny